OpenStudy (anonymous):

how do you set up the equation when you are trying to Find the sum of a arithmetic series? (how would i plug it in a graphing calculator?)

OpenStudy (amistre64):

the basic construct of the sum of a series is to add the first and last terms of the series together in such a way that it doubles the actual sum, for example: 1 + 3 + 5 + 7 + 9 9 + 7 + 5 + 3 + 1 ----------------- 10+10+10+10+10 = (9+1), 5 times 5(9+1) = 50, which is double what we need, so divide it in half 1 + 3 + 5 + 7 + 9 = 25

OpenStudy (amistre64):

\[\S_n=\frac{n}{2}(a_1+a_n)\]

OpenStudy (amistre64):

thats an odd looking S .....

OpenStudy (anonymous):

thank you :)